# Latent Functional Maps

### v2

@@ 4.3 LFMs as a similarity measure
Once computed, the functional map $\mathbf{C}$ can serve as a measure of similarity between spaces. The reason is that for isometric transformations between manifolds, the functional map is volume preserving (see Thm 5.1 in [34]), and this is manifested in orthogonal $\mathbf{C}$. By defining the inner product between functions $h_{1},h_{2}\in\mathcal{F}({M})$ as $\langle h_{1},h_{2}\rangle=\int_{\mathcal{M}}h_{1}(x)h_{2}(x)\mu(x)$, it holds that $\langle h_{1},h_{2}\rangle=\langle\hat{h}_{1},\hat{h}_{2}\rangle$ when the map preserves the local area, where $\hat{h}$ denotes the functional representation of $h$. In other words, when the transformation between the two manifolds is an isometry, the matrix $\mathbf{C}^{T}\mathbf{C}$ will be diagonal. By measuring the ratio between the norm of the off-diagonal elements of $\mathbf{C}^{T}\mathbf{C}$ and the norm of its diagonal elements, we can define a measure of similarity $sim(X,Y)=1-\frac{||\text{off}((\mathbf{C}^{T}\mathbf{C})||_{F}}{||\text{diag}(\mathbf{C}^{T}\mathbf{C})||_{F}}$. Furthermore, this quantity is interpretable; the first eigenvector of $\mathbf{C}^{T}\mathbf{C}$ can act as a signal to localize the area of the target manifold where the map has higher distortion [33].
@@ 4.4 Transfering information with LFM
The functional map computed between two latent spaces can be utilized in various ways to transfer information from one space to the other. In this paper, we focus on two methods: (i) Expressing arbitrary points in the latent space as distance function on the graph and transferring them through the functional domain; (ii) Obtaining a point-to-point correspondence between the representational spaces from the LFM, starting from none to few known pairs, and leverage off-the-shelf methods to learn a transformation between the spaces. Additional strategies could be explored in future work.
Space of Functional Coefficients
The space of functional coefficients offers an alternative representation for points in the latent space $\mathcal{X}$. Using the equation $\hat{f}_{G}=\mathbf{\Phi}_{G}^{T}f_{G}$, any function $f_{G}\in\mathcal{F}(G,\mathbb{R})$ can be uniquely represented by its functional coefficients $\hat{f}_{G}$. We leverage this property to represent any point $x\in\mathcal{X}$ as a distance function $f_{d}\in\mathcal{F}(G,\mathbb{R})$ from the set of points $X_{G}$, which correspond to the nodes of the graph $G$. The functional map $\mathbf{C}$ between two latent spaces $\mathcal{X}$ and $\mathcal{Y}$ aligns their functional representations, enabling the transfer of any function from the first space to the second. This functional alignment can be used similarly to the method proposed by [29] to establish a "relative" space where the representational spaces $\mathcal{X}$ and $\mathcal{Y}$ are aligned.
Extending spaces correspondences
As explained in Section 3, the functional map $\mathbf{C}$ represents the bijection $T$ in a functional form. [34] demonstrated that this bijection can be retrieved as a point-to-point map by finding the nearest neighbor for each row of $\mathbf{\Phi}_{G_{Y}}\mathbf{C}$ in $\mathbf{\Phi}_{G_{X}}$. This process can be efficiently implemented using algorithms such as kd-tree. Given a few correspondences (anchors) between the two spaces $\mathcal{X}$ and $\mathcal{Y}$, we can extend these correspondences to the entire set of nodes $X$ and $Y$. This extended set of anchors can then be used to determine a transformation between the latent spaces, as described by [22]. In the following section, we demonstrate that by using a small number of anchors (typically $\leq 50$), we can retrieve optimal transformations that facilitate near-perfect stitching and retrieval.
@@ 5.1 Analysis
We demonstrate the benefits of using latent functional maps for comparing distinct representational spaces, using the similarity metric defined in Section4.3
@@
In order to validate experimentally if LFMs can serve as a good measure of similarity between distinct representational spaces, we run the same sanity check as in [14]. We train 10 CNN models (the architecture is depicted in Appendix B.1) on the CIFAR-10 dataset [15], changing the initialization seed. We compare their inner representations at each layer, excluding the logits and plot them as a similarity matrix, comparing with Central Kernel Alignment (CKA) measure [14] and Canonical Correlation Analysis (CCA) [11, 37]. We then measure the accuracy of identifying corresponding layers across models and report the results comparing with CKA and CCA as baselines. For CCA, we apply average pooling on the spatial dimensions to the embeddings of the internal layers, making it more stable numerically and boosting the results for this baseline compared to what was observed in [14].
@@ 5.2 Zero-shot stitching
We test the use of the latent functional map in the task of zero-shot stitching, as defined in [29], to combine independent encoders and decoders (e.g., classifiers, generators) without subsequent training or fine-tuning.
@@
(b) Fine labelling
Figure 4: Stitching on CIFAR100.Accuracy performance in stitching between image encoders trained on CIFAR100, comparing orthogonal transformation (Ortho) and LFM+Ortho at varying anchor counts. Also shown is LFM+Ortho (Labels), which uses the dataset labels instead of anchors. Results are presented for (a) coarse and (b) fine-grained labeling, with mean accuracy values reported on each box.
@@
We consider four pre-trained image encoders (see Appendix B.2 for details) and stitch their latent spaces to perform classification using a Support Vector Machine (SVM) on two different labelings of CIFAR100 [15]: coarse and fine-grained. To evaluate the effectiveness of integrating the functional map, we extend the correspondences to determine an orthogonal transformation [22] between the latent spaces. For each encoder, we compute a graph of 3,000 points with 300 neighbors per node. We optimize the problem in Equation 2 using the first 50 eigenvectors of the graph Laplacian and consider two different descriptors: the distance functions defined from the anchors (LFM+Ortho) and the labels (LFM+Ortho (Labels)). For each dataset class, the latter provides an indicator function with 1 if the point belongs to the class and 0 otherwise. This descriptor type does not require any anchor as input, representing a pioneering example of stitching requiring no additional information beyond the dataset.
@@
Figure 4 presents the accuracy results for all possible combinations of encoder stitching. The addition of the latent functional map (LFM+Ortho) shows higher performance with a low number of anchors in both labelings of CIFAR100. Even without any anchors, the label descriptors (LFM+Ortho (Labels)) provide the best performance for the latent functional map framework in both labelings. Computing the orthogonal transformation directly from the anchors (Ortho) proves to have comparable performance only with 500 anchors, where the performance of LFM is limited by the number of eigenvectors used.
@@ 5.3 Retrieval
For this experiment, we construct k-NN graphs with a neighborhood size of 300 and compute the functional map using the first 50 eigenvectors. We evaluate the methods’ performance using the Mean Reciprocal Rank (MRR), as detailed in Appendix B.4. Our functional map methods are compared with the method proposed by [29] (Relatives) and the orthogonal transformation method proposed by [22] (Ortho).
@@
Figure 5 shows the performance of these methods as the number of anchors increases. The numerical results are detailed in Table 2 in the Appendix. The functional map significantly improves performance with just 5 anchors, achieving an MRR of over 0.8. As the number of anchors increases, the performance of competing methods improves but still falls short of FMAP+Transform at 300 anchors, which reaches an MRR of 0.99. Interestingly, the performance of the functional map methods does not improve beyond 5 anchors, suggesting that this number of anchors is sufficient to achieve an optimal functional map between the spaces. In Appendix C.1, we further analyze how the results improve as the number of eigenvectors used to compute the functional map increases. Notably, the MRR score drastically increases to over 0.6 with more than 25 eigenvectors.
These results confirm that the latent functional map is a valuable tool in settings with little knowledge about correspondences. It significantly enhances retrieval performance with a minimal number of anchors, making it an efficient approach for aligning latent spaces. Moreover, its performance can be improved using a higher number of eigenvectors.
@@ 6 Conclusions
In this paper, we explored the application of latent functional maps (LFM) to better understand and utilize the internal representations of neural networks. Our approach leverages the principles of spectral geometry to model and analyze the relationships between different latent spaces, providing a robust framework for comparing and aligning representations across various models. By extending the functional map framework to high-dimensional latent spaces, we offer a versatile method for comparing distinct representational spaces, finding correspondences between them in both unsupervised and weakly supervised settings (few or no anchors) and transferring information across different models.
@@
The performance of LFM can be sensitive to the number of eigenvectors used, as seen in some of our experiments. Finding the optimal number of eigenvectors without extensive experimentation remains an open problem. The current framework may not handle very complex transformations between latent spaces as effectively as simpler linear transformations. Expanding the framework to manage more complex transformations is an area for future exploration. Future research can further explore the potential of LFM in other domains and tasks, and modalities. The versatility of the functional representation can be further explored to define new ad-hoc constraints and regularizers for different settings. In particular, ts performance in fully unsupervised settings (with no anchors) is still an area requiring further research and improvement.
In conclusion, the latent functional map framework offers a novel and effective approach to understanding and utilizing neural network representations, promising significant advancements in both theoretical and practical aspects of machine learning.
Appendix A Latent Functional Map
A.1 Additional Regularizers
In Equation 2, we improve the computation of the functional map by incorporating two additional regularizers: Laplacian commutativity and descriptor operator commutativity. Both regularizers exploit the preservation of linear functional operators $\mathbf{S}^{G}:\mathcal{F}(G,\mathbb{R})\rightarrow\mathcal{F}(G,\mathbb{R})$, enforcing that the functional map $\mathbf{C}$ commutes with these operators: $||\mathbf{S}^{G}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}||=0$.
@@ A.1 Additional Regularizers
$\rho_{\mathcal{L}}(\mathbf{C})=||\mathbf{\Lambda}_{G_{Y}}\mathbf{C}-\mathbf{C}\mathbf{\Lambda}_{G_{X}}||^{2}$ (3)
@@
$\rho_{f}(\mathbf{C})=\sum_{i}||\mathbf{S}^{G_{Y}}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}||^{2}$ (4)
where $\mathbf{S}^{G}_{i}=\mathbf{\Phi}_{G}^{T}\text{Diag}(f^{G}_{i})\mathbf{\Phi}_{G}$ are the descriptor operators.

### v3

@@ 3.3 LFMs as a similarity measure
Once computed, the functional map $\mathbf{C}$ can serve as a measure of similarity between spaces. The reason is that for isometric transformations between manifolds, the functional map is volume preserving (see Thm 5.1 in [44]), and this is manifested in orthogonal $\mathbf{C}$. By defining the inner product between functions $h_{1},h_{2}\in\mathcal{F}({M})$ as $\langle h_{1},h_{2}\rangle=\int_{\mathcal{M}}h_{1}(x)h_{2}(x)\mu(x)$, it holds that $\langle h_{1},h_{2}\rangle=\langle\hat{h}_{1},\hat{h}_{2}\rangle$ when the map preserves the local area, where $\hat{h}$ denotes the functional representation of $h$. In other words, when the transformation between the two manifolds is an isometry, the matrix $\mathbf{C}^{T}\mathbf{C}$ will be diagonal. By measuring the ratio between the norm of the off-diagonal elements of $\mathbf{C}^{T}\mathbf{C}$ and the norm of the full matrix, we can define a measure of similarity $sim(X,Y)=1-\frac{||\text{off}((\mathbf{C}^{T}\mathbf{C})||_{F}^{2}}{||\mathbf{C}^{T}\mathbf{C}||_{F}^{2}}$. In Appendix A.4 we prove that this similiarity measure is a proper distance on the space of functional maps, allowing to compare measurements from collections of spaces. Furthermore, this quantity is interpretable; the first eigenvector of $\mathbf{C}^{T}\mathbf{C}$ can act as a signal to localize the area of the target manifold where the map has higher distortion [43], as we show in the experiment in Figure 3(a).
@@ 3.4 Transfering information with LFM
The map $\mathbf{C}$ computed between two latent spaces can be utilized in various ways to transfer information from one space to the other. In this paper, we focus on two methods: (i) Expressing arbitrary points in the latent space as distance function on the graph and transferring them through the functional domain; (ii) Obtaining a point-to-point correspondence between the representational spaces from the LFM (see the first step in Appendix A.3) , starting from none to few known pairs, and leverage off-the-shelf methods to learn a transformation between the spaces. Additional strategies could be explored in future work.
Space of Functional Coefficients.
The space of functional (or spectral) coefficients offers an alternative representation for points in the latent space $\mathcal{X}$. Using the equation $\hat{f}_{G}=\mathbf{\Phi}_{G}^{T}f_{G}$, any function $f_{G}\in\mathcal{F}(G,\mathbb{R})$ can be uniquely represented by its functional coefficients $\hat{f}_{G}$. We leverage this property to represent any point $x\in\mathcal{X}$ as a distance function $f_{d}\in\mathcal{F}(G,\mathbb{R})$ from the set of points $X_{G}$, which correspond to the nodes of the graph $G$. The functional map $\mathbf{C}$ between two latent spaces $\mathcal{X}$ and $\mathcal{Y}$ aligns their functional representations, enabling the transfer of any function from the first space to the second. This functional alignment can be used similarly to the method proposed by [39] to establish a shared space where the representational spaces $\mathcal{X}$ and $\mathcal{Y}$ are aligned.
Finding correspondences.
As explained in Section 3, the functional map $\mathbf{C}$ represents the bijection $T$ in a functional form. [44] demonstrated that this bijection can be retrieved as a point-to-point map by finding the nearest neighbor for each row of $\mathbf{\Phi}_{G_{Y}}\mathbf{C}$ in $\mathbf{\Phi}_{G_{X}}$. This process can be efficiently implemented and scaled up using kd-tree structures or approximation strategies [21, 20]. Starting from an empty set or few known correspondences (anchors) between the two spaces $\mathcal{X}$ and $\mathcal{Y}$, we can extend the correspondence to the entire set of nodes $X$ and $Y$. This extended set of anchors can then be used to fit an arbitrary transformation between the latent spaces, for example an orthogonal mapping [32]. In our experiments, we demonstrate that by using a very small number of anchors (typically $\leq 50$), we can retrieve optimal transformations that facilitate near-perfect stitching and retrieval tasks.
@@ 4 Experiments
In this section, we present a series of experiments designed to evaluate the effectiveness of the Latent Functional Map (LFM) framework. We explore its application across various tasks, including similarity evaluation, stitching, retrieval performance, and robustness to perturbations in latent space. By comparing LFM to existing methods under different conditions, we aim to demonstrate its versatility and superior performance in aligning and analyzing neural network representations. Additional ablations and qualitative visualizations on the choice of distance metric to construct the graph and descriptors selection are reported in the Appendix C.
@@ 4.1 Analysis
We demonstrate the benefits of using latent functional maps for comparing distinct representational spaces, using the similarity metric $sim(X,Y)$ defined in Section 3.3
@@
In order to validate experimentally if LFMs can serve as a good measure of similarity between distinct representational spaces, we run the same sanity check as [23]. We train 10 CNN models (the architecture is depicted in Appendix B.1) on the CIFAR-10 dataset [24], changing the initialization seed. We compare their inner representations at each layer, excluding the logits and plot them as a similarity matrix, comparing with Central Kernel Alignment (CKA) measure [23] and Canonical Correlation Analysis (CCA) [18, 47]. We then measure the accuracy of identifying corresponding layers across models and report the results comparing with CKA and CCA as baselines. For CCA, we apply average pooling on the spatial dimensions to the embeddings of the internal layers, making it more stable numerically and boosting the results for this baseline compared to what was observed in [23].
@@ 4.2 Zero-shot stitching
In latent communication tasks, a common challenge is combining an encoder that embeds data with a decoder specialized in a downstream task, such as classification or reconstruction—this process is often referred to as stitching. Previous works, like [29] and [2], have used trainable linear projections, known as stitching layers, to enable the swapping of different network components. However, [39] introduced the concept of zero-shot stitching, where neural components can be marged without any additional training procedure. It’s important to note that while [39] still required to trained a decoder module once for processing relative representations before stitching could be performed, our method is fully zero-shot, with no need for additional training. In the following experiments, stitching is conducted without any training or fine-tuning of the encoder or decoder, adhering strictly to a zero-shot fashion.
@@
(b) CIFAR100 fine
[…]
(c) ImageNet
[…]
(d) CUB
[…]
(e) MNIST
[…]
(f) AgNews
Figure 4: Stitching results. Accuracy performance in stitching between image encoders trained on 6 datasets, comparing orthogonal transformation (Ortho) and LFM+Ortho at varying anchor counts. Also shown is LFM+Ortho (Labels), which uses the dataset labels instead of anchors. Results are presented with mean accuracy values reported on each box.
[…]
(a) MNIST
[…]
(b) FashionMNIST
[…]
(c) CIFAR-10
Figure 5: Qualitative results on stitching. We combine the encoder decoder modules of two convolutional autoencoder (E1 D1, E2 D2) using three different approaches: no transformation (Absol.), the orthogonal transformation from [32] (Ortho), the orthogonal transformation augmented with LFM (LFM).
@@
We consider four pre-trained image encoders (see Appendix B.2 for details) and stitch their latent spaces to perform classification using a Support Vector Machine (SVM) on five different datasets: CIFAR100 [24] with coarse and fine-grained labelling, ImageNet [11], Caltech-UCSD Birds-200-2011 (CUB) [58], MNIST [12], AgNews [64]. These datasets were chosen to encompass both large-scale, complex datasets like ImageNet and a variety of modalities, including text and images. To evaluate the effectiveness of integrating the functional map, we extend the correspondences to determine an orthogonal transformation [32] between the latent spaces. For each encoder, we compute a graph of 3,000 points with 300 neighbors per node. We optimize the problem in Equation 2 using the first 50 eigenvectors of the graph Laplacian and consider two different descriptors: the distance functions defined from the anchors (LFM+Ortho) and the labels (LFM+Ortho (Labels)). For each dataset class, the latter provides an indicator function with 1 if the point belongs to the class and 0 otherwise. This descriptor does not require any anchor as input, representing a pioneering example of stitching requiring no additional information beyond the dataset.
@@
Figure 4 shows the accuracy results for all possible combinations of encoder stitching. The latent functional map (LFM+Ortho) consistently outperforms other methods across all datasets, particularly with a low number of anchors. Even without anchors, using label descriptors (LFM+Ortho (Labels)) achieves strong performance across different labeling schemes. The orthogonal transformation computed directly from the anchors (Ortho) only matches LFM’s performance when more than 50 anchors are used, at which point LFM’s effectiveness is constrained by the number of eigenvectors. Comparing the CIFAR100 coarse and fine-grained labeling, the fine-grained task is more complex due to the larger number of classes, making the estimation of the map more challenging. This complexity is especially evident when aligning self-supervised vision models with classification-based ones (e.g., DINO vs. ViT), where the training strategies differs.
@@
4.2.1 Qualitative example
In Figure 5, we present qualitative experiments on the MNIST, FashionMNIST [63], and CIFAR-10 datasets, focusing on visualizing the reconstructions of stitched autoencoders. For these experiments, we trained convolutional autoencoders using two different seeds and stitched their encoder and decoder modules together. As a baseline, we used no transformation (Absol.), where the latent representation from the first encoder is directly input into the second decoder. We then compared this baseline with the orthogonal transformation from [32] (Ortho) and its LFM-augmented version (LFM). For both our method and [32], we used 10, 10 and 50 correspondences for the MNIST, FashionMNIST, and CIFAR-10 datasets, respectively. Our method consistently produced superior reconstruction quality compared to both the baseline and the method from [32].
@@ 4.3 Retrieval
For this experiment, we construct k-NN graphs with a neighborhood size of 300 and compute the functional map using the first 50 eigenvectors. We evaluate the methods’ performance using the Mean Reciprocal Rank (MRR), as detailed in Appendix B.4. Our functional map methods are compared with the method proposed by [39] (Relatives) and the orthogonal transformation method proposed by [32] (Ortho). We choose to fit the same orthogonal transformation on top of the correspondence found by LFM to make the comparison the fairest possible, although in principle any off-the-shelf methods could be used to estimate the transformation once the new correspondence is found. We illustrate this versatility in Table 3 of the Appendix, where LFM is combined with other methods in the same task.
@@
Figure 6 shows the performance of these methods as the number of anchors increases. Numerical results are detailed in Table 3 in the Appendix. The functional map significantly improves performance with just 5 anchors, achieving an MRR of over 0.8. As the number of anchors increases, the performance of competing methods improves but still falls short of FMAP+Transform at 300 anchors, which reaches an MRR of 0.99. Interestingly, the performance of the functional map methods does not improve beyond 5 anchors, suggesting that this number of anchors is sufficient to achieve an optimal functional map between the spaces. These findings are further supported by Table 3 in the Appendix, where LFM consistently achieves superior MRR scores, particularly with fewer anchors, outperforming other baselines such as [29]. In Appendix C.3, we analyze how the results improve as the number of eigenvectors used to compute the functional map increases. Notably, the MRR score drastically increases to over 0.6 with more than 25 eigenvectors.
These results confirm that the latent functional map is a valuable tool in settings with little knowledge about correspondences. It significantly enhances retrieval performance with a minimal number of anchors, making it an efficient approach for aligning latent spaces. Moreover, its performance can be improved using a higher number of eigenvectors. In particular, our framework comprises at the same time (i) an interpretable similarity measure, (ii) a way to find correspondences, and (iii) solving for an explicit mapping between different spaces in a single framework, differently from [39] and [32] which attempt to solve just the latter implicitly.
@@ 5 Conclusions
In this paper, we explored the application of latent functional maps (LFM) to model and analyze the relationships between different latent spaces. Our approach leverages the principles of spectral geometry, providing a robust framework for comparing and aligning representations across various models and settings. By extending the functional map framework to high-dimensional latent spaces, we offer a versatile method for comparing distinct representational spaces, finding correspondences between them in both unsupervised and weakly supervised settings (few or no anchors) and transferring information across different models.
@@
The performance of LFM can be sensitive to the number of eigenvectors used, as observed in some of our experiments (see Appendix C.3). Finding the optimal number of eigenvectors without extensive experimentation remains an open problem. The current framework assumes a correspondence between the domains, and is not directly adapted in order to deal with partial mappings, where just a subset of one domain can be mapped to other. Extending the results in [48] to our neural representation setting, is a promising way to overcome the full correspondence assumption. Future research can further explore the potential of LFM in other domains and tasks, and modalities. The versatility of the functional representation can be further explored to define new ad-hoc constraints and regularizers for different settings. In particular, ts performance in fully unsupervised settings, while promising is still an area requiring further research and improvement. In conclusion, the Latent Functional Map framework offers a novel and effective approach to understanding and utilizing neural network representations, promising significant advancements in both theoretical and practical aspects of machine learning.
Acknowledgments and Disclosure of Funding
MF is supported by the MSCA IST-Bridge fellowship which has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101034413. ER and VM are supported by the PNRR MUR project PE0000013-FAIR. MP is supported by the Sapienza grant "Predicting and Explaining Clinical Trial Outcomes", prot. RG12218166FA3F13.
Appendix A Latent Functional Maps
A.1 Building the graph
In the following section we give more details about how to construct the graph in the latent space in order to approximate the manifold domain from a sample estimate.
Throughout the paper, we assume the eigenvalues (and corresponding eigenvectors) are sorted in non-descending order $0=\Lambda_{1}\leq\Lambda_{2}\leq\dots\leq\Lambda_{n}$. One may consider a subset of eigenvectors, namely those associated with the $k$ smallest eigenvalues, to compactly approximate a graph signal, employing techniques akin to Fourier analysis. As demonstrated in multiple recent works [53, 7], the eigenvalues and eigenvectors of the graph Laplacian associated with a k-NN graph approximate the weighted Laplace-Beltrami operator, placing us in a setting similar to the original one of [44]. In particular in [7] it has been derived optimal bounds for the choices of $k$, bounds based on the dimensionality of the underlying manifold and the number of sampled points. In our experiments we choose our parameter $k$ according to these estimates. For an estimate of manifold dimensionality we use the latent space dimensionality, which correspond to an upper bound to it. In practice, we typically set $k=300$ with a graph in the order of few thousands nodes.
A.2 Additional Regularizers
In Equation 2, we improve the computation of the functional map by incorporating two additional regularizers: Laplacian commutativity and descriptor operator commutativity. Both regularizers exploit the preservation of linear functional operators $\mathbf{S}^{G}:\mathcal{F}(G,\mathbb{R})\rightarrow\mathcal{F}(G,\mathbb{R})$, enforcing that the functional map $\mathbf{C}$ commutes with these operators: $\|\mathbf{S}^{G}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}\|_{F}=0$.
@@ A.2 Additional Regularizers
$\rho_{\mathcal{L}}(\mathbf{C})=\|\mathbf{\Lambda}_{G_{Y}}\mathbf{C}-\mathbf{C}\mathbf{\Lambda}_{G_{X}}\|_{F}^{2}$ (3)
@@
$\rho_{f}(\mathbf{C})=\sum_{i}\|\mathbf{S}^{G_{Y}}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}\|_{F}^{2}$ (4)
where $\mathbf{S}^{G}_{i}=\mathbf{\Phi}_{G}^{T}(f^{G}_{i})\mathbf{\Phi}_{G}$ are the descriptor operators.
A.3 Spectral refinement
Once we have solved the optimization problem defined in Equation 2, we refine the resulting functional map $\mathbf{C}$ using the spectral upsampling algorithm proposed by [34]. We start by considering the matrix $\mathbf{C}$ as the submatrix of size $k\times k$ of the current LFM estimate. Then we apply iteratively the following two steps:
  1. 1.
[…]
  2. 2.
    Set $\mathbf{C}_{k+1}={\mathbf{\Phi}_{G_{X}}^{:,k}}^{T}\mathbf{\Pi}\mathbf{\Phi}_{G_{Y}}^{:,k}$, where $\mathbf{\Phi}^{:,k}$ correspond to the $k$-th column of the matrix $\mathbf{\Phi}$.
A.4 LFM similarity properties
In this section we will demonstrate that the LFM similarity property $sim(X,Y)=1-\frac{\|\text{off}((\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}{\|(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}$, defined in Section 3.3 is a metric, as it satisfies, positiveness, simmetry and triangle inequality. Namely, for a pairs of spaces $X,Y$:
$sim(X,Y)\geq 0$ (5)
$sim(X,Y)=sim(Y,X)$ (6)
$sim(X,Z)\leq sim(X,Y)+sim(Y,Z)$ (7)
Since $\|\text{diag}(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}+\|\text{off}(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}=\|(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}$ we will consider $sim(X,Y)=sim(X,Y)=\frac{\|\text{diag}(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}{\|(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}$ for simplicity. The former property is trivial as the ration between two norms is positive. For proving simmetry we have that:
$sim(X,Y)=\frac{\|\text{diag}(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}{\|(\mathbf{C}^{T}\mathbf{C})\|_{F}^{2}}$ (8)
$sim(X,Y)=\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}{\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}$ (9)
for the numerator $\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}$ we trivially have $diag(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})=diag(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X})$. For the denominator:
Since :$\text{Since :}\|\mathbf{\Phi}\|_{F}=\sqrt{\sum_{i,j}|\mathbf{\Phi}_{i,j}|^{2}}=\sqrt{Tr(\mathbf{\Phi}^{T}\mathbf{\Phi})}$ (10)
We have that:
$\|\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y}\|_{F}^{2}\stackrel{{\scriptstyle?}}{{=}}\|\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X}\|_{F}^{2}$ (11)
$Tr(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})^{T}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\stackrel{{\scriptstyle?}}{{=}}Tr(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X})^{T}(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X})$ (12)
$Tr(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X}\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})=Tr(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y}\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{X})$ (13)
where the last equality is verified for the Trace operator being invariant under cyclic permutations.
For proving triangle inequality we will assume that the matrices of eigenvectors are full rank (i.e. all eigenvectors are considered). We have to prove the following:
$1-\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}$ $\leq 2-(\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}{\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}+\frac{\|\text{diag}((\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{\|(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}})$ (14)
$\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}$ $\geq(\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}{\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}+\frac{\|\text{diag}((\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{\|(\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}})-1$ (15)
Using the fact that the Frobenius norm is invariant under multiplication by orthogonal matrices the denominator of each term in Eq14 becomes: $\|(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}=\|\mathbf{\Phi}_{X}\|_{F}^{2}=Tr[\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{X}]=Tr[Id]=N$. Then:
$\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{N}\geq(\frac{\|\text{diag}(\mathbf{\Phi}_{X}^{T}\mathbf{\Phi}_{Y})\|_{F}^{2}}{N}+\frac{\|\text{diag}((\mathbf{\Phi}_{Y}^{T}\mathbf{\Phi}_{Z})\|_{F}^{2}}{N})-1$ (16)
Since $\|\text{diag}(\mathbf{X}_{X}^{T}\mathbf{Y})\|_{F}^{2}\leq\|(\mathbf{X}_{X}^{T}\mathbf{Y})\|_{F}^{2}$ for any $\mathbf{X},\mathbf{Y}$ then the RHS of Eq. 16 will always be smaller than the LHS.